how to write the slope-intercept form of 4x-2y=3 through given points of (2,1)into parallel and perpendicular forms?
Answer by mananth(16946)

PARALLEL


4 x -2 y = 3
Find the slope of this line
make y the subject
-2 y = -4 x + 3
Divide by -2
y = 2 x -1 1/2
Compare this equation with y=mx+b
slope m = 2
The slope of a line parallel to the above line will be the same
The slope of the required line will be 2
m= 2 ,point ( 2 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = 4 + b
b= -3
m= 2
Plug value of the slope and b in y = mx +b
The required equation is y = 2 x -3
Perpendicular
4 x + -2 y = 3
Find the slope of this line

-2 y = -4 x + 3
Divide by -2
y = 2 x + -1 1/2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 2

The slope of a line perpendicular to the above line will be the negative reciprocal - 1/2
Because m1*m2 =-1
The slope of the required line will be - 1/2

m= - 1/2 ,point ( 2 , 1 )
Find b by plugging the values of m & the point in
y=mx+b
1 = -1 + b
b= 2
m= - 1/2
The required equation is y = - 1/ 2 x 2
m.ananth@hotmail.ca